OpenStudy (anonymous):

Given ratio=3x first term=1 and infinite sum=4/9 determine the value of x

OpenStudy (anonymous):

3x or 3 ???

OpenStudy (anonymous):

so... 1/1-3x= 4/9 ===>9=4-12x

OpenStudy (anonymous):

how did you get 9=4-12x?

OpenStudy (anonymous):

sum of an finite gp is a/1-r

OpenStudy (anonymous):

I'm still not sure how you got to that answer

OpenStudy (bibby):